# A Unified Understanding and Evaluation of Steering Methods

### v1

@@ Abstract
Steering methods provide a practical approach to controlling large language models by applying steering vectors to intermediate activations, guiding outputs toward desired behaviors while avoiding retraining. Despite their growing importance, the field lacks a unified understanding and consistent evaluation across tasks and datasets, hindering progress. This paper introduces a unified framework for analyzing and evaluating steering methods, formalizing their core principles and offering theoretical insights into their effectiveness. Through comprehensive empirical evaluations on multiple-choice and open-ended text generation tasks, we validate these insights, identifying key factors that influence performance and demonstrating the superiority of certain methods. Our work bridges theoretical and practical perspectives, offering actionable guidance for advancing the design, optimization, and deployment of steering methods in LLMs.
Machine Learning, ICML
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across diverse tasks, from language understanding to creative generation (Touvron et al., 2023; Anthropic, 2023; OpenAI, 2023). However, their inherent generality also poses challenges in controlling their behavior to meet specific objectives or align with desired behaviors. Very recently, steering has emerged as a promising approach to modulating the behavior of LLMs, enabling fine-grained control without requiring additional training or significant architectural modifications (Li et al., 2024; Rimsky et al., 2024; Zou et al., 2023). Steering methods operate by applying constant shifts or steering vectors to intermediate model activations, nudging the model away from generating undesirable responses and toward desired generations.
@@
To address the problem, this paper provides a unified framework for understanding and evaluating steering methods, and identifying the principles that govern their effectiveness or ineffectiveness. This framework captures the core principles underlying established steering methods, such as CAA (Rimsky et al., 2024), RepE (Zou et al., 2023), and ITI (Li et al., 2024), enabling direct comparisons across techniques and fostering a deeper understanding of their mechanisms. As illustrated in Figure 1(a), these methods typically operate on a dataset of contrastive pairs, with positive generations that align with the target behavior and negative generations that do not. For example, CAA learns a steering vector by computing the mean of differences between embeddings of positive and negative examples. RepE extracts the first principal component of the differences between embeddings of positive and negative examples. ITI uses a binary classifier trained on the embeddings from positive vs. negative generations and uses the learned parameters as the steering vector. This unified framework provides critical insights into steering design, enabling better method selection and optimization. Our theoretical analysis in Section 3 shows that for steering negative examples towards positive examples, the mean of differences is the optimal steering vector (see formally in Theorem 3.1).
Going beyond theoretical understanding, our framework provides a comprehensive evaluation of steering approaches, isolating the effects of how steering vectors are learned while carefully controlling for other design factors (Section 4). Our evaluation encompasses a diverse range of tasks, including both multiple-choice questions and open-ended text generation. Our results reveal that the mean of differences (Rimsky et al., 2024) consistently outperforms PCA-based and classifier-based approaches across most tasks, aligning well with our theoretical insight. Furthermore, visualization of embeddings shows that PCA-based approach (Zou et al., 2023) may struggle to perform effectively, especially in scenarios when the positive and negative embeddings vary along a direction that is nearly orthogonal to the steering vector. These findings validate our theoretical insights, demonstrating their practical relevance.
@@
  • We present a unified mathematical characterization of established steering methods, and provide theoretical results demonstrating the optimality of certain steering methods while explaining the limitations of others.
@@ 2 Preliminaries on Steering Methods
Figure 1: (Top) Illustration of the contrastive pair from the myopic-reward dataset (Perez et al., 2022). (Bottom) Resulting steering vectors from different methods where there exists a perfect steering vector $\mathbf{v}^{*}$. See Section 3.1 for details.
@@
To elucidate strategies available for steering LLMs, we describe methods for learning steering vectors based on several established frameworks: Contrastive Activation Addition (CAA) (Rimsky et al., 2024), Representation Engineering (RepE) (Zou et al., 2023), and Inference-Time Intervention (ITI) (Li et al., 2024). Most methods for learning steering vectors are based upon a contrastive dataset, which consists of positive generations that align with the target behavior and negative generations which do not align with the target behavior. For example, Rimsky et al. (2024) employed prompt pairs consisting of multiple-choice questions with answer letters appended at the end. While both prompts in each pair present the same question, they end with different answers; the “positive” prompt concludes with the letter corresponding to the desired behavior, and the “negative” prompt ends with a letter indicative of the contrary behavior. We refer to this pair of examples as a contrastive pair and provide an example in Figure 1(a). We denote a dataset of contrastive pairs as $\left\{(\mathbf{x}_{i},\mathbf{y}_{i}^{(+)},\mathbf{y}_{i}^{(-)})\right\}_{i=1}^{N}$ with $N$ examples, where $\mathbf{x}_{i}$ is the prompt, $\mathbf{y}_{i}^{(+)}$ is the positive response, and $\mathbf{y}_{i}^{(-)}$ is the negative response.
@@
TopPC gives the first principal component normalized to have unit length, and $\text{Classify}(\cdot)$ gives a learned vector for classifying embeddings from positive responses versus embeddings from negative responses and is normalized to have length equal to the standard deviation along that direction, following the practice in Li et al. (2024).
@@ 3. PCA of Embeddings
4. Classifier (Li et al., 2024)
In the case of a binary classifier trained with cross-entropy loss, the resulting steering vector may approximate $\mathbf{v}^{*}$ in direction. During the learning process, the first gradient step starting from zero or near-zero initialization gives a direction close to $\mathbf{v}^{*}$, which we show in Appendix A. However, subsequent steps steps may deviate significantly, especially when the steering vector is derived from a limited number of samples. Even if the direction may be similar to $\mathbf{v}^{*}$ due to choices such as the learning rate and number of steps, the scale of the vector may not be accurate. While one can normalize the vector using the standard deviation along its direction, as performed in ITI (Li et al., 2024), this does not guarantee a vector with length equivalent to $\mathbf{v}^{*}$. Figure 1(d) illustrates the resulting vector, where it is evident that although the learned direction is similar to $\mathbf{v}^{*}$, the magnitude of the steering vector is substantially shorter than $\mathbf{v}^{*}$.
@@ 3.2 Formal Guarantee
Given any joint distribution of contrastive pairs $\Pi_{+,-}$, such that the marginal distributions over positive and negative embeddings have finite means, and the objective given in (8), then the steering vector that minimizes the objective is the mean of differences:
@@
Theorem 3.1 provides a foundational result for understanding steering methods by formalizing the relationship between contrastive datasets and the optimal steering vector. A key implication of this theorem is that the mean difference (Rimsky et al., 2024) serves as the optimal steering vector. The theorem also highlights the data-dependent nature of steering methods. Since the optimal vector is derived from the mean difference across the joint distribution of contrastive pairs, the quality of the dataset can impact the performance of the steering method. For example, steering is inherently constrained by the separability of positive and negative embeddings within the given dataset. Additionally, while the mean difference approach minimizes the average error across the entire distribution, it does not guarantee optimal performance for every individual sample. Such cases can arise from outliers or data points that deviate significantly from the assumptions underlying the contrastive dataset.
Dataset CAA RepE Top PC ITI
(Rimsky et al., 2024) (Zou et al., 2023) (Li et al., 2024)
@@ 4 Empirical Verification
We build on the experimental setup of Rimsky et al. (2024), performing comparative analyses of various methods for learning steering vectors. Our datasets consist of multiple alignment-relevant behaviors sourced from Anthropic’s evaluation datasets (Perez et al., 2022) and datasets generated by (Rimsky et al., 2024), including (1) Coordination with Other AIs , (2) Corrigibility, (3) One Box Tendency, (4) Refusal, (5) Self-awareness, (6) Survival Instinct, and (7) Wealth Seeking. For each dataset, we comprehensively test all the steering methods presented in Section 2 and evaluate the resulting generation in both multiple-question and open-ended generation scenarios. Following the original setup in Rimsky et al. (2024), we use Llama-7B-Chat (Touvron et al., 2023). For multiple-choice questions, we measure the performance using the average probability of the correct answer (APC) and the average accuracy (ACC). For open-ended generation, we consider the model’s response to a free-form question with the steering vector applied, and use GPT-4 to score responses based on how well they align with the target behavior.
@@ 4.1 Multiple-Choice QA
Table 1 summarizes the steering performance on different QA tasks. Each dataset contains multiple-choice questions with two answer options that demonstrate either the behavior of interest or its opposite, as illustrated in Figure 1(a). We observe that for the majority of behaviors tested, the mean difference method in Rimsky et al. (2024) consistently outperforms other methods by a large margin. The empirical observation matches well with our Theorem 3.1.
It’s worth noting that we have made a rigorous effort to ensure a fair comparison across all methods. In particular, we perform steering for all methods at the same layer and use embeddings extracted from the same location within the transformer architecture. The optimal layer (layer 13) and location (i.e., residual stream) are identified based on our detailed ablation study, as discussed further in Section 4.3. These optimal configurations are also consistent with findings in prior work. Moreover, to achieve the best possible performance for each method, we use a validation set to choose an optimal multiplier for the steering vector from the set of $\{0.5,1,1.5,2,2.5,3\}$, and report the performance on a held-out test set. For the classifier-based method (Li et al., 2024), we learn the steering vector using the Binary Cross-Entropy loss and apply gradient descent for 1000 steps to ensure convergence.
@@
Self-Awareness 14.43 12 32.62 36 32.88 36 17.6 16
Survival Instincts 22.05 18 33.7 30 33.58 28 34.48 26
@@
Visualization of steering vectors. In Figure 2, we provide a visualization of the embeddings to demonstrate how the intuitions from the illustrative example translate to real-world datasets. The visualizations are based on the embeddings for the behavior of Corrigibility from the Anthropic dataset. For each steering method, we project the embeddings onto the learned steering vector along the $x$-axis, and to the top principal component of the orthogonal subspace along the $y$-axis. Ideally, the positive embeddings in blue and the negative embeddings in red should be distinctly separated along the steering direction, or the horizontal axis. However, while the mean of differences and classifier methods achieve this separation, the PCA-based methods do not. For PCA-based method employed in RepE (Zou et al., 2023), the positive and negative embeddings vary along a direction that is nearly orthogonal to the steering vector. This behavior results from the mismatch between the direction with the most variance and the direction corresponding to the desired behavior. Accordingly, the PCA-based methods have the lowest performance (cf. Table 1).
@@ 4.2 Open-Ended Generation
We further test the different steering methods on free-form answers to open-ended questions, as shown in Table 3. Following the setup of Rimsky et al. (2024), we adapt held-out multiple choice questions into open-ended prompts by providing only the initial question without answer options. We form open-ended contrastive pairs by appending the correct and incorrect responses directly after the question, and we use the embeddings at the last step to learn the steering vector. Similar to the multiple-choice task, we apply the steering vector at the residual stream of the optimal layer to test set generation questions. We use GPT-4 to rate the responses on a scale of 1-10 based on how well the response aligns with the target behavior of the dataset. In order to evaluate the effect of the steering vector, we consider the difference in scores after applying the steering vector with a multiplier of +1/1+1/-1+ 1 / - 1. We see that similarly to the multiple-choice setting, the mean difference vector performs best for most of the datasets. For the remaining datasets, we find that the mean difference vector has the second best performance following the classifier. This suggests that the findings and theory derived from steering vectors learned from contrastive pairs generalize to open-ended generation. We find similar results when considering the raw scores which we provide in Appendix C. The prompt template for GPT-4 evaluation is attached in Appendix B.
Dataset CAA RepE Top PC ITI
$\Delta$Score $\Delta$Score $\Delta$Score $\Delta$Score
Coordination 0.22 0.04 0.08 -0.1
Corrigible 2.28 -0.16 0.5 7.58
@@
Refusal 0.4 -0.04 0.08 0
Self-Awareness 0.76 0 0 0
@@
Wealth Seeking 0.44 0.12 0 2.52
Table 3: Open-ended behavioral evaluation. For all methods, steering is performed at layer 13. $\Delta$Score indicates the difference between positive and negative steering in ratings given by GPT-4 based on how well the response aligns with the behavior being steered. Higher numbers are better.

### v2

@@ Abstract
Latent space steering methods provide a practical approach to controlling large language models by applying steering vectors to intermediate activations, guiding outputs toward desired behaviors while avoiding retraining. Despite their growing importance, the field lacks a unified understanding and consistent evaluation across tasks and datasets, hindering progress. This paper introduces a unified framework for analyzing and evaluating steering methods, formalizing their core principles and offering theoretical insights into their effectiveness. Through comprehensive empirical evaluations on multiple-choice and open-ended text generation tasks, we validate these insights, identifying key factors that influence performance and demonstrating the superiority of certain methods. Our work bridges theoretical and practical perspectives, offering actionable guidance for advancing the design, optimization, and deployment of latent space steering methods in LLMs.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across diverse tasks, from language understanding to creative generation [49, 1, 35]. However, their inherent generality also poses challenges in controlling their behavior to meet specific objectives or align with desired behaviors. Very recently, latent space steering has emerged as a promising approach to modulating the behavior of LLMs, enabling fine-grained control without requiring additional training directly on target outputs or significant architectural modifications [28, 41, 57]. Latent space steering methods operate by applying constant shifts or steering vectors to intermediate model activations, nudging the model away from generating undesirable responses and toward desired generations. These steering vectors are learned in an unsupervised manner or with the addition of binary labels based on some desired property.
@@
To address the problem, this paper provides a unified framework for understanding and evaluating steering methods, and identifying the principles that govern their effectiveness or ineffectiveness. This framework captures the core principles underlying established steering methods, such as CAA [41], RepE [57], and ITI [28], enabling direct comparisons across techniques and fostering a deeper understanding of their mechanisms. As illustrated in Figure 1(a), these methods typically operate on a dataset of contrastive pairs, with positive generations that align with the target behavior and negative generations that do not. For example, CAA learns a steering vector by computing the mean of differences between embeddings of positive and negative examples. RepE extracts the first principal component of the differences between embeddings of positive and negative examples. ITI uses a binary classifier trained on the embeddings from positive vs. negative generations and uses the learned parameters as the steering vector. This unified framework provides critical insights into steering design, enabling better method selection and optimization. Our theoretical analysis in Section 3 shows that for steering negative examples towards positive examples, the mean of differences is the optimal steering vector (see formally in Theorem 3.1).
Going beyond theoretical understanding, our framework provides a comprehensive evaluation of steering approaches, isolating the effects of how steering vectors are learned while carefully controlling for other design factors (Section 4). Our evaluation encompasses a diverse range of tasks, including both multiple-choice questions and open-ended text generation. Our results reveal that the mean of differences [41] consistently outperforms PCA-based and classifier-based approaches across most tasks, aligning well with our theoretical insight. Furthermore, visualization of embeddings shows that PCA-based approach [57] may struggle to perform effectively, especially in scenarios when the positive and negative embeddings vary along a direction that is nearly orthogonal to the steering vector. These findings validate our theoretical insights, demonstrating their practical relevance.
@@
  • We present a unified mathematical characterization of established latent space steering methods, and provide theoretical results demonstrating the optimality of certain steering methods while explaining the limitations of others.
@@ 2 Preliminaries on Steering Methods
Figure 1: (Top) Illustration of the contrastive pair from the myopic-reward dataset [38]. (Bottom) Resulting steering vectors from different methods where there exists a perfect steering vector $\mathbf{v}^{*}$. See Section 3.1 for details.
@@
To elucidate strategies available for steering LLMs, we describe methods for learning steering vectors based on several established frameworks: Contrastive Activation Addition (CAA) [41], Representation Engineering (RepE) [57], and Inference-Time Intervention (ITI) [28]. Most methods for learning steering vectors are based upon a contrastive dataset, which consists of positive generations that align with the target behavior and negative generations which do not align with the target behavior. For example, Rimsky et al. [41] employed prompt pairs consisting of multiple-choice questions with answer letters appended at the end. While both prompts in each pair present the same question, they end with different answers; the “positive” prompt concludes with the letter corresponding to the desired behavior, and the “negative” prompt ends with a letter indicative of the contrary behavior. We refer to this pair of examples as a contrastive pair and provide an example in Figure 1(a). We denote a dataset of contrastive pairs as $\left\{(\mathbf{x}_{i},\mathbf{y}_{i}^{(+)},\mathbf{y}_{i}^{(-)})\right\}_{i=1}^{N}$ with $N$ examples, where $\mathbf{x}_{i}$ is the prompt, $\mathbf{y}_{i}^{(+)}$ is the positive response, and $\mathbf{y}_{i}^{(-)}$ is the negative response.
@@
TopPC gives the principal component corresponding to the largest singular value normalized to have unit length, and $\text{Classify}(\cdot)$ gives a learned vector for classifying embeddings from positive responses versus embeddings from negative responses and is normalized to have length equal to the standard deviation along that direction, following the practice in Li et al. [28].
@@ 3. PCA of Embeddings
4. Classifier [28]
In the case of a binary classifier trained with cross-entropy loss, the resulting steering vector may approximate $\mathbf{v}^{*}$ in direction. During the learning process, the first gradient step starting from zero or near-zero initialization gives a direction close to $\mathbf{v}^{*}$, which we show in Appendix A. However, subsequent steps steps may deviate significantly, especially when the steering vector is derived from a limited number of samples. Even if the direction may be similar to $\mathbf{v}^{*}$ due to choices such as the learning rate and number of steps, the scale of the vector may not be accurate. While one can normalize the vector using the standard deviation along its direction, as performed in ITI [28], this does not guarantee a vector with length equivalent to $\mathbf{v}^{*}$. Figure 1(d) illustrates the resulting vector, where it is evident that although the learned direction is similar to $\mathbf{v}^{*}$, the magnitude of the steering vector is substantially shorter than $\mathbf{v}^{*}$.
@@ 3.2 Formal Guarantee
Given any joint distribution of contrastive pairs $\Pi_{+,-}$, such that the marginal distributions over positive and negative embeddings have finite means, and the objective given in equation 8, then the steering vector that minimizes the objective is the mean of differences:
@@
Theorem 3.1 provides a foundational result for understanding steering methods by formalizing the relationship between contrastive datasets and the optimal steering vector. A key implication of this theorem is that the mean difference [41] serves as the optimal steering vector. The theorem also highlights the data-dependent nature of steering methods. Since the optimal vector is derived from the mean difference across the joint distribution of contrastive pairs, the quality of the dataset can impact the performance of the steering method. For example, steering is inherently constrained by the separability of positive and negative embeddings within the given dataset. Additionally, while the mean difference approach minimizes the average error across the entire distribution, it does not guarantee optimal performance for every individual sample. Such cases can arise from outliers or data points that deviate significantly from the assumptions underlying the contrastive dataset.
Mitigating Limitations
As our theoretical results imply, steering vectors are limited in their ability to adapt precisely to a broad range of samples. We discuss some recent methods that attempt to steer in a more controlled and fine-grained manner and some potential alternatives to be explored. Recent works such as [25, 45, 51, 8, 11] consider learning multiple vectors specific to a cluster of embeddings or learning a condition such as a linear classifier. Other works such as [4, 43, 42] consider learning a linear function rather than a steering vector. Other potential directions include learning non-linear but relatively simple models such as a kernel-based model or a higher-order function.
Dataset MoD PoD PoE CoE
@@ 4 Empirical Verification
We build on the experimental setup of Rimsky et al. [41], performing comparative analyses of various methods for learning steering vectors. Our datasets consist of multiple alignment-relevant behaviors sourced from Anthropic’s evaluation datasets [38] and datasets generated by [41], including (1) Coordination with Other AIs , (2) Corrigibility, (3) One Box Tendency, (4) Refusal, (5) Self-awareness, (6) Survival Instinct, and (7) Wealth Seeking. For each dataset, we comprehensively test all the steering methods presented in Section 2 and evaluate the resulting generation in both multiple-question and open-ended generation scenarios. Following the original setup in Rimsky et al. [41], we use Llama-2-7b-Chat [49]. For multiple-choice questions, we measure the performance using the average probability of the correct answer (APC) and the average accuracy (ACC). For open-ended generation, we consider the model’s response to a free-form question with the steering vector applied, and use GPT-4 to score responses based on how well they align with the target behavior. We provide additional results on Mistral-7B-Instruct-v0.3 [22] and Llama-3.1-8B-Instruct [17] in Appendix B.
@@ 4.1 Multiple-Choice QA
Table 1 summarizes the steering performance on different QA tasks. Each dataset contains multiple-choice questions with two answer options that demonstrate either the behavior of interest or its opposite, as illustrated in Figure 1(a). We observe that for the majority of behaviors tested, the mean difference method in Rimsky et al. [41] consistently outperforms other methods by a large margin. The empirical observation matches well with our Theorem 3.1.
It’s worth noting that we have made a rigorous effort to ensure a fair comparison across all methods. In particular, we perform steering for all methods at the same layer and use embeddings extracted from the same location within the transformer architecture. The optimal layer (layer 13) and location (i.e., residual stream) are identified based on our detailed ablation study, as discussed further in Section 4.3. These optimal configurations are also consistent with findings in prior work. Moreover, to achieve the best possible performance for each method, we use a validation set to choose an optimal multiplier for the steering vector from the set of $\{0.5,1,1.5,2,2.5,3\}$, and report the performance on a held-out test set. For the classifier-based method [28], we learn the steering vector using the Binary Cross-Entropy loss and apply gradient descent for 1000 steps to ensure convergence.
@@
Self-Awareness 14.43 12 32.62 36 32.88 36 17.60 16
Survival Instincts 22.05 18 33.70 30 33.58 28 34.48 26
@@
Visualization of steering vectors. In Figure 2, we provide a visualization of the embeddings to demonstrate how the intuitions from the illustrative example translate to real-world datasets. The visualizations are based on the embeddings for the behavior of Corrigibility from the Anthropic dataset. For each steering method, we project the embeddings onto the learned steering vector along the $x$-axis, and to the top principal component of the orthogonal subspace along the $y$-axis. Ideally, the positive embeddings in blue and the negative embeddings in red should be distinctly separated along the steering direction, or the horizontal axis. However, while the mean of differences and classifier methods achieve this separation, the PCA-based methods do not. For PCA-based method employed in RepE [57], the positive and negative embeddings vary along a direction that is nearly orthogonal to the steering vector. This behavior results from the mismatch between the direction with the most variance and the direction corresponding to the desired behavior. Accordingly, the PCA-based methods have the lowest performance (cf. Table 1).
@@ 4.2 Open-Ended Generation
We further test the different steering methods on free-form answers to open-ended questions, as shown in Table 3. Following the setup of Rimsky et al. [41], we adapt held-out multiple choice questions into open-ended prompts by providing only the initial question without answer options. We form open-ended contrastive pairs by appending the correct and incorrect responses directly after the question, and we use the embeddings at the last step to learn the steering vector. Similar to the multiple-choice task, we apply the steering vector at the residual stream of the optimal layer to test set generation questions. We use GPT-4 to rate the responses on a scale of 1-10 based on how well the response aligns with the target behavior of the dataset. In order to evaluate the effect of the steering vector, we consider the difference in scores after applying the steering vector with a multiplier of +1/1+1/-1. We see that similarly to the multiple-choice setting, the mean difference vector performs best for most of the datasets. For the remaining datasets, we find that the mean difference vector has the second best performance following the classifier. This suggests that the findings and theory derived from steering vectors learned from contrastive pairs generalize to open-ended generation. We find similar results when considering the raw scores which we provide in Tables 4 and 5. The prompt template for GPT-4 evaluation is attached in Appendix C.
Dataset MoD PoD PoE CoE
$\DeltaScore Δ\DeltaScore$ $\DeltaScore Δ\DeltaScore$
Coordination 0.22 0.04 0.08 -0.10
Corrigible 2.28 -0.16 0.50 7.58
@@
Refusal 0.40 -0.04 0.08 0.00
Self-Awareness 0.76 0.00 0.00 0.00
@@
Wealth Seeking 0.44 0.12 0.00 2.52
[…]
Dataset MoD PoD PoE CoE
Score Score Score Score
Coordination 0.50 0.24 0.16 0.36
Corrigible 7.04 5.16 5.66 7.82
One Box 8.24 7.50 7.60 7.39
Refusal 10.0 4.54 4.42 10.0
Self-Awareness 9.16 9.00 9.00 9.00
Survival 5.30 5.40 5.34 5.98
Wealth Seeking 8.90 8.94 8.82 9.38
Table 4: Open-ended behavioral evaluation raw positive scores. For all methods, steering is performed at layer 13. The scores are ratings given by GPT-4 based on how well the response aligns with the behavior being steered. For positive steering, higher numbers are better.
Dataset MoD PoD PoE CoE
Score Score Score Score
Coordination 0.28 0.20 0.08 0.46
Corrigible 4.76 5.32 5.16 0.24
One Box 5.58 7.96 6.67 7.30
Refusal 9.60 4.58 4.34 10.0
Self-Awareness 8.40 9.00 9.00 9.00
Survival 4.41 5.26 5.24 4.52
Wealth Seeking 8.46 8.82 8.82 6.86
Table 5: Open-ended behavioral evaluation raw negative scores. For all methods, steering is performed at layer 13. The scores are ratings given by GPT-4 based on how well the response aligns with the behavior being steered. For negative steering, lower numbers are better.